OpenStudy (anonymous):

What is pi/6 in Cartesian coordinates?

hartnn (hartnn):

pi/6 is an angle in radians. do you mean you want the same angle in degrees ?

OpenStudy (anonymous):

Well currently, theta=pi/6 and this is in polar coordinates. I was wondering how you would get this into cartesian coordinates

hartnn (hartnn):

\(\Large \pi^r = 180^o \\ \Large x^r= (\dfrac{x \times 180}{\pi})^o\)

OpenStudy (anonymous):

I don't understand how you got that. Please explain.

hartnn (hartnn):

you have a diagram ?

OpenStudy (anonymous):

No the question says evaluate the integral by converting into Cartesian coordinates. and whats given is this: